OpenStudy (anonymous):

Determine the value of the co-efficient "B" F(s)=7s^2+4s+2/(s+1)(s^2+s+5)=(A/s+1)+(Bs+C/s^2+s+5

OpenStudy (anonymous):

no one seems to want to do this because it is a pain

OpenStudy (anonymous):

you have to write \[A(x^2+x+5)+(Bx+C)(x+1)=7x^2+4x+2\] then multiply out. then equate like coefficients. then find them

OpenStudy (anonymous):

you get \[Ax^2+Ax+5A+Bx^2+Bx+Cx+C=7x^2+4x+2\]

OpenStudy (anonymous):

giving 3 equations \[A+B=7\] \[A+B+C=4\] \[5A+C=2\]

OpenStudy (anonymous):

i guess we can subtract the first from the second and get C = -3

OpenStudy (anonymous):

replacing C by -3 in the third gives \[5A-3=2\] so A=1 and finally replacing A by 1 equation 1 gives B = 6 what a pita

OpenStudy (anonymous):

Thank you , you are the MASTER !!!

OpenStudy (anonymous):

yes now i am a master. graduated from superhero

OpenStudy (anonymous):

now you can explain these fourier series to me